Consider the function:

10x2 + 2x + 2.5 = f(s)

Obtain:

1.) the transfer function X(s)/F(s) treating f(t) as a unit step function

2.) the transfer function X(s)/F(s) treating f(t) as a unit impulse function

Plot both responses using Matlab and explain the differences
Answer #1

Program:

%Note: This program will work only if you have control system tool box in
%your matlab


%f=10*x^2 + 2*x + 2.5; given function
%H=1/f=1/(10*x^2 + 2*x + 2.5)% transfer function
sys = tf(1,[10 2 2.5]); % obtaining transfer function
subplot(211); step(sys); % treating as unit step function
subplot(212); impulse(sys); % treating as unit impulse function